### v1

# Formatting instructions for NeurIPS 2018

Abstract
Retraining machine learning models remains an important task for real-world machine learning model deployment. Existing methods focus largely on greedy approaches to find the best-performing model without considering the stability of trained model structures across different retraining evolutions. In this study, we develop a mixed integer optimization algorithm that holistically considers the problem of retraining machine learning models across different data batch updates. Our method focuses on retaining consistent analytical insights—which is important to model interpretability, ease of implementation, and fostering trust with users—by using custom-defined distance metrics that can be directly incorporated into the optimization problem. Importantly, our method shows stronger stability than greedily trained models with a small, controllable sacrifice in model performance in a real-world production case study. Finally, important analytical insights, as demonstrated using SHAP feature importance, are shown to be consistent across retraining iterations.
1 Introduction
As more machine learning models are adopted in real-world production pipelines, there is an increasing demand for interpretability to obtain key analytical insights which can build trust in AI-driven decision support systems [1, 2, 3]. Such model understanding enables developers to evaluate a model’s reliability, which is crucial for high-stake decision spaces such as politics and healthcare [4, 5, 6].
However, when a machine learning model is updated as new data is incorporated into training, analytical insights often diverge significantly from previous observations [7]. As users develop trust in models over time based on their past performance and model alignment with personal expertise, when insights suddenly shift due to new data update, it can lead to skepticism and hesitation for adoption .
Consider a predictive model used to estimate a patient’s mortality in a hospital: if the model’s insights drastically change with each new iteration, physicians may find it challenging to explain these variations to their patients and may even create legal difficulties if incidents of error arise. Patients rely on doctors’ recommendations, and if the predictions keep fluctuating, it can challenge their confidence in the medical decision-making process [4].
There are multiple avenues to address challenges faced by retraining in the literature, via continual learning [8], online learning [9], as well as many others [10, 11, 12, 13]. However, existing methods focus largely on retaining prior performance on the retrained models in the face of data distributional shifts, or conceptual shifts, and thus do not put an emphasis on the fact that analytical insights: such as the decision path of a prediction, or important features selected by a model, be stable across different retraining iterations. In this work, we demonstrate a direct incorporation of analytical insights into training more stable models and ensure that the output models provide consistent analytical outcomes.
Additionally, retraining should be distinguished from online learning, which aims at continuously updating data and more importantly, incrementally updating the machine learning model instead of retraining the model from scratch with each data batch update (batch learning). One advantage of batch learning is its smaller consumption of computational resources and faster retraining speed since it is done in a periodical instead of a continuous fashion. Furthermore, such a training strategy allows for a more global solution, as retraining involves using the entire database, whereas online learning, often to ensure faster speed, incrementally incorporates only a portion of the original data, and provides more greedy solutions [14].
Recently, another body of literature focusing on model structure stability has been investigated in the operations research and computer science community. Various methods, such as mixed integer optimization, were used in [15, 16] to achieve sequentially stable regression and classification tree models. Other works including [17] also developed methods that produce stable feature interactions in random forest-based models. Our work follows and extends the work of [16] by developing a unifying framework that learns stable structure across multiple different modeling techniques, including regression, classification trees, and gradient-boosted trees.
1.1 Contributions
We summarize the main contributions of this paper: (i) we develop a general framework aimed at stabilizing machine learning model structures and formulate it as a mixed-integer optimization-based algorithm for retraining machine learning models that improve existing greedy approaches. (ii) we demonstrate a significant boost in model stability with small, computational efficiency, and controlled sacrifice of performance. (iii) we show its applicability in an existing production pipeline used in a large US hospital for mortality prediction.
The rest of the paper is organized as follows. In Section 2, we outline the problem setting and formulation of the retraining setting, in Section 3, we present numerical experimental results on a real-world case study, in 4, we discuss our limitations and conclude our study.
2 A Methodology for Retraining Machine Learning Models
In this section, we introduce the problem setting of the retraining problem, its formulation as a mixed integer problem, and finally discuss the adaptive training strategies.
2.1 Problem Setting
Consider an observational dataset composed of triplets $(\boldsymbol{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\boldsymbol{x_{i}}\in\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathbb{R}$ is the target variable, and $t_{i}$ is the time at which the data point $i$ was obtained. The dataset is accumulated over time, resulting in a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$. The dataset available at time $T_{b},\ b\in[B]:=\{1,\ldots,B\},$ is denoted by $D_{b}$ and is defined as:
$D_{b}=\{(x_{i},y_{i}):t_{i}\leq T_{b},i\in[N]\}.$
Thus, we have that $D_{1}\subseteq D_{2}\subseteq\ldots\subseteq D_{B}$. We assume that $B\geq 2$ and that, at model training time (denoted by $t$), all $B$ batches are available, i.e., $t\geq T_{B}$. In Section 2.6, we discuss how to adaptively retrain the model using batches obtained after time $t$.
2.2 Problem Statement
The goal is to find models $f_{1},f_{2},\ldots,f_{B}$ with $f_{b}:\mathbb{R}^{p}\mapsto\mathbb{R}$ for $b\in[B]$ such that the following two requirements are satisfied:
Minimizing Predictive Error: The first requirement is to ensure that each model $f_{b}$ in the sequence $f_{1},\ldots,f_{B}$ has a low out-of-sample predictive error. Formally, denoting by $\mathcal{D}_{b}$ the distribution of the training data in $D_{b}$, we seek to minimize the expected error for data points $(\boldsymbol{x},y)\sim\mathcal{D}_{b}$:
$\min_{f_{1},\ldots,f_{B}}\sum_{b=1}^{B}\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}_{b}}[\ell(f_{b}(\boldsymbol{x}),y)],$ (1)
where the loss function $\ell(\cdot,\cdot)$ quantifies the prediction error when the predicted value is $f_{b}(\boldsymbol{x})$ and the true value is $y$.
Ensuring Model Stability: The second requirement is to ensure that consecutive model structures do not vary too significantly, maintaining stability over time. Formally, we define a stability metric $d(f_{i},f_{j})$ that measures the distance (change) between models $i$ and $j$. Then, the objective is to minimize this metric for each pair of consecutive models:
$\min_{f_{1},\ldots,f_{B}}\sum_{b=1}^{B-1}d(f_{b},f_{b+1}).$ (2)
We remark that the stability metric could be based on parameters (in parametric models), structure (in tree-based models), or even predictions.
In practice, we are primarily interested in model $f_{B}$ as this model has been trained on the full training set and is the one that will be used in production. Nevertheless, we choose to train the full sequence $f_{1},\ldots,f_{B}$ of $B$ models for the following reasons. First, by training the full sequence, we favor models that change smoothly upon the addition of new batches of data. Second, forming the full sequence allows us to observe changing patterns, which is crucial for interpretability purposes. Therefore, our approach relies on imposing stability by process.
2.3 A General Formulation
Recall our assumption that, at training time, we have already received $B\geq 2$ batches of data. (If this is not the case, we artificially split the training data into $B$ batches.) The requirements stated in Equations (1) and (2) naturally yield the following bi-objective optimization problem:
$\min_{f_{1},\ldots,f_{B}}\left(\lambda\sum_{b=1}^{B}L(f_{b},D_{b})+(1-\lambda)\sum_{b=1}^{B-1}d(f_{b},f_{b+1})\right),$ (3)
where $L(f,D)$ measures the predictive error of model $f$ on data $D$ and $\lambda$ is a weighting parameter that balances the two objectives. However, in practice, we often want to explicitly control “the price of stability” with respect to predictive power. Formally, given a suboptimality tolerance parameter $\alpha\geq 0$, the predictive error of each model $f_{b}$ should not be worse than the best model for batch $D_{b}$ by more than a multiplicative factor of $\alpha$. This motivates the following constrained formulation:
$\begin{split}\min_{f_{1},\ldots,f_{B}}&\quad\sum_{b=1}^{B-1}d(f_{b},f_{b+1})\\ \text{s.t.}&\quad L(f_{b},D_{b})\leq(1+\alpha)\cdot\min_{f}L(f,D_{b})\quad\forall b\in[B].\end{split}$ (4)
Depending on the type of the model used, solving Problem (4) can be challenging, as we need to consider either an infinite number (i.e. neural network) or an exponentially growing number (with respective to feature size, i.e. decision tree) of possible functions $f_{b}$. This is particularly true for models that are commonly used in high-stakes healthcare applications, e.g., decision trees and variants thereof. Thus, we next develop a restricted formulation that can tractably accommodate any type of machine learning model.
2.4 A Tractable Restricted Formulation
For each batch $b\in[B]$, we train a finite set of candidate models, denoted as $\mathcal{F}_{b}$. Then, the problem reduces to selecting one model from each set of pre-computed models to form a sequence. Let $f_{b,j}$ be the $j$-th model in $\mathcal{F}_{b}$, and introduce binary variables $z_{b,j}$ such that $z_{b,j}=1$ if model $f_{b,j}$ is selected, and 0 otherwise. This yields the following MIO formulation:
$\begin{split}\min_{\begin{subarray}{c}f_{b,j}\in\mathcal{F}_{b}\\ z_{b,j}\in\{0,1\}\end{subarray}}&\quad\sum_{b=1}^{B-1}\sum_{j=1}^{|\mathcal{F}_{b}|}\sum_{k=1}^{|\mathcal{F}_{b+1}|}d(f_{b,j},f_{b+1,k})\cdot z_{b,j}\cdot z_{b+1,k}\\ \text{s.t.}&\quad\sum_{j=1}^{|\mathcal{F}_{b}|}L(f_{b,j},D_{b})\cdot z_{b,j}\leq(1+\alpha)\cdot\min_{f\in\mathcal{F}_{b}}L(f,D_{b})\quad\forall b\in[B]\\ &\quad\sum_{j=1}^{|\mathcal{F}_{b}|}z_{b,j}=1\quad\forall b\in[B].\end{split}$ (5)
The following theorem allows us to solve the restricted problem of selecting a slowly varying sequence of models (Problem (5)) in polynomial time:
Theorem 2.1 (Reduction of Model Selection to Shortest Path Problem).
Given a sequence of batches $\{D_{b}\}_{b=1}^{B}$ with associated sets of pre-computed candidate models $\{\mathcal{F}_{b}\}_{b=1}^{B}$, and a stability metric $d(f_{b,j},f_{b+1,k})$ for any pair of models $f_{b,j}$ in $\mathcal{F}_{b}$ and $f_{b+1,k}$ in $\mathcal{F}_{b+1}$, Problem (5) can be reduced to a shortest path problem in a directed graph.
Proof.
Proof of Theorem 2.1 We construct the reduction as follows:
  1. 1.
    Graph Construction: Construct a directed graph $G=(V,E)$ where each node $v_{b,j}$ corresponds to a model $f_{b,j}$ from each model set $\mathcal{F}_{b}$ for each batch $b$. Include additional nodes $v_{\text{source}}$ and $v_{\text{sink}}$. Build directed edges from each node in $\mathcal{F}_{b}$ to each node in $\mathcal{F}_{b+1}$, and from $v_{\text{source}}$ to each node in $\mathcal{F}_{1}$ and from each node in $\mathcal{F}_{B}$ to $v_{\text{sink}}$.
  2. 2.
    Edge Weight Assignment: Assign a weight of $d(f_{b,j},f_{b+1,k})$ to each edge $(v_{b,j},v_{b+1,k})$. Assign a weight of zero to edges from $v_{\text{source}}$ to $\mathcal{F}_{1}$ and from $\mathcal{F}_{B}$ to $v_{\text{sink}}$.
  3. 3.
    Node Filtering: For each batch $b$, include only those nodes $v_{b,j}$ corresponding to models $f_{b,j}$ that satisfy $L(f_{b,j},D_{b})\leq(1+\alpha)\cdot\min_{f}L(f,D_{b})$.
  4. 4.
    Shortest Path Problem: Finding a sequence of models that minimizes the total stability metric while satisfying the performance constraints is equivalent to finding the shortest path from $v_{\text{source}}$ to $v_{\text{sink}}$ in graph $G$.
  5. 5.
    Polynomial-Time Solvability: Since the shortest path problem in a directed graph with non-negative weights can be solved in polynomial time (e.g., using Dijkstra’s algorithm), the model selection problem can also be solved in polynomial time.
2.5 Model Distance Calculation
To compute the distance $d(f_{i},f_{j})$ between models $f_{i}$ and $f_{j}$, we consider two primary approaches: structural distances and feature importance-based distances.
Structural Distances: This approach measures the distance by directly comparing the models’ structures. We provide the following concrete examples:
  • For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\beta}^{i}-\boldsymbol{\beta}^{j}\|_{2}^{2}$, where $\boldsymbol{\beta}^{i}$ and $\boldsymbol{\beta}^{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively.
  • For decision tree-based models, we use the decision tree distance metric proposed by [16] that relies on optimally matching the paths in the two trees $\mathbb{T}_{1}$ and $\mathbb{T}_{2}$, denoted as $\mathcal{P}(\mathbb{T}_{1})$ and $\mathcal{P}(\mathbb{T}_{2})$. The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account a path weight $w(p)$, defined as the differences in feature ranges in the two paths as well as the class labels associated with them. Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem. This distance is defined as an integer linear optimization problem, where given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{1}$ is matched with path $q$ in $\left.\mathbb{T}_{2}\right)$ and $x_{p}=$ $\mathbb{I}$ (path $p$ in $\mathbb{T}_{1}$ is left unmatched), we formulate it as below:
    $d\left(\mathbb{T}_{1},\mathbb{T}_{2}\right)=\min_{\boldsymbol{x}}$ $\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}\sum_{q\in\mathcal{P}\left(\mathbb{T}_{2}\right)}d(p,q)x_{pq}+\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}w(p)x_{p}$
    s.t. $\sum_{q\in\mathcal{P}\left(\mathbb{T}_{2}\right)}x_{pq}+x_{p}=1,\quad\forall p\in\mathcal{P}\left(\mathbb{T}_{1}\right)$
    $\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}x_{pq}=1,\quad\forall q\in\mathcal{P}\left(\mathbb{T}_{2}\right)$
    $x_{pq}\in\{0,1\},\quad x_{p}\in\{0,1\},\quad\forall p\in\mathcal{P}\left(\mathbb{T}_{1}\right),\quad\forall q\in\mathcal{P}\left(\mathbb{T}_{2}\right)$
Feature Importance-Based Distances: This approach measures the distance by comparing the models based on their respective feature importances. In particular, denoting by $\boldsymbol{\phi}^{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\phi}^{f_{i}}-\boldsymbol{\phi}^{f_{j}}\|_{2}^{2}$. The advantage of using feature importance-based distances is its versatility to be applied to both interpretable (decision-tree and logistic regression), as well as black-box (gradient boosted tree) models. Concretely, we consider the following two approaches to define feature importance measures:
  • For decision tree-based models, we use the Gini importance, which quantifies how much a feature contributes to homogenizing the target variable within the subsets created by the splits. Given $v\in V$ as a child node within the tree, $I(j,v)$ as the Gini impurity reduction at node $v$ by feature $j$ and $I(v)$ as the Gini impurity reduction at node $v$ regardless of feature, feature importance of feature $j$, $FI_{j}$, is computed as its normalized total reduction of the Gini impurity:
    $FI_{j}=\frac{\sum_{v\in V}I(j,v)}{\sum_{v\in V}I(v)}.$
  • For general black-box models, we propose using SHAP (SHapley Additive exPlanations) values [18], which provide a unified approach to feature importance across various well-known black-box models. Given feature $j$, its SHAP feature importance is computed as the average marginal SHAP value contribution, $\phi_{j}^{(i)}$, across all possible samples $i$:
    $FI_{j}=\frac{1}{n}\sum_{i=1}^{n}|\phi_{j}^{(i)}|.$
2.6 Adaptive Model Retraining
We now discuss how to apply this framework in the context of adaptive model retraining—when new batches become available and we wish to update an existing slowly varying sequence to incorporate the new data. More specifically, we have some observational dataset $D_{B}$ already available with an existing slowly varying sequence $f_{1},f_{2},\dots,f_{B}$. We are now given an additional batch of data which we append to the existing dataset to create $D_{B+1}$. We want to obtain a model $f_{B+1}$ that remains close in distance to the current model, $f_{B}$.
To do so, we suggest the construction of an independent slowly varying sequence $f^{\prime}_{1},f^{\prime}_{2},\dots,f^{\prime}_{B+1}$ created on the updated dataset, $D_{B+1}$, and deploying $f^{\prime}_{B+1}$. Unlike the perhaps more obvious choice of greedily choosing a new model $f^{\mathrm{greedy}}_{B+1}$ trained on $D_{B+1}$ that is close to $f_{B}$, this approach achieves stability by process and generalizes to longer deployment horizons with multiple model updates and can also be used when re-training after receiving multiple data updates—e.g., constructing $f^{\prime}_{B+m}$ when the existing model is $f_{m}$.
3 Numerical Experiments
3.1 Case Study
Dataset.   Daily patient data from the largest hospital system in Connecticut are collected and patient features ranging from demographics, patient status, clinical measurements, laboratory results, diagnoses, orders, procedures, and notes are extracted. Patient records are retrospectively identified from January 1st, 2018 to May 1st, 2022, overseeing 168,815 patients, and 865,954 patient-day records. Using the available data, we focus on the prediction of patient mortality risk, defined as the probability of each patient expiring or going to hospice at the end of the hospital stay.
Data Preprocessing.   In the existing production pipeline, models are trained on a 6-month basis, although the retraining window is scheduled to be reduced to 3-4 months in 2024. There are no obvious data shifts for this case study. To simulate the process of receiving data updates, we divide the 52 months into multiple sub-intervals. We follow traditional machine learning training procedures by dividing the dataset into training, validation, and testing sets. All patient records admitted from September 2021 to December 2022 (16 months) are allocated as the out-of-sample testing set. The dataset and problem setting discussed in this case study can be found in more detail in [19].
Research questions.   To assess the performance of our slowly varying methodology, we evaluate both the intra-sequence stability of the generated sequences—i.e., the pairwise distance of adjacent models chosen in a slowly varying sequence—as well as the inter-sequence stability of the sequences—i.e., the distance between the final models of independent, slowly varying sequences. The intra-sequence stabilityresults serve to illuminate the distance-accuracy trade-off that our methodology hopes to optimize and the inter-sequence results to analyze the stability of the adaptive model retraining strategy. A secondary research question is the ideal update frequency (or interval length) for the SVML methodology, we begin the case study by empirically evaluating multiple different interval lengths, before fixing an interval length of 3 months for the inter- and intra-sequence experiments.
Experimental Methodology.   For each of the three research questions, we obtain experimental results as follows. Having defined an interval length, the number of batches $B$ is fixed. For each batch $b\in B$, we define the available training data $D_{b}$ as the first $b$ intervals and reserve the $b+1$th interval as a validation set. We generate a set of pre-computed candidate models, $\mathcal{F}_{b}$, for each batch, bootstrapping to ensure diversity in the candidate models. Then, for accuracy tolerances [0.1, 0.05, 0.02,0.01], we generate a slowly varying sequence according to the methodology outlined in 2.4. To evaluate the performance of these sequences, we also generate a “greedy sequence” constructed of greedily selected, best-performing models to simulate a sequence of models (evaluated on the validation set) without regard to stability. We denote by $f_{b}$ the batch $b$ model chosen by the slowly varying sequence and denote by $f^{\ast}_{b}$ the batch $b$ model chosen by the greedy sequence. All AUC results are reported on the 16-month testing set put aside at the start of model training.
Experimental Software.   We implement all algorithms in Python 3.7. We solve the optimization models using Gurobi 9.5. All experiments were performed on a standard Intel(R) Xeon(R) CPU E5-2690 @ 2.90GHz running CentOS release 7.
3.2 Results
Optimal interval length.   Though in practice this real-world dataset is updated on a 6-month basis, given that we have access to 52 months of data, we investigate empirically the update frequency that leads to the best performance of our slowly varying methodology. We chose the XGBoost model as the canonical model type for the experiment since this is the model in use in production. We performed four trials, 1-, 3-, 6-, and 9-month intervals, following the experimental methodology discussed with 25 pre-computed candidate models for each batch with results in Fig 1. 3-month interval was chosen as the optimal re-training frequency as it does not exhibit strong unstable behavior compared to 1-month, and can retain high AUC ranging from 0.85-0.9 with improvements in distance from [0.3-0.37] to [0.275-0.21] compared to 6- or 9-month intervals.
Figure 1: The AUC and distance between adjacent XGBoost modelsfor each interval length.
Intra-sequence stability.   Having selected 3-month as the optimal update frequency, we turn towards a more in-depth analysis of the distance-accuracy trade-off implicit in the slowly varying methodology. We conduct this experiment for three classes of models: logistic regression, CART decision trees, and XGBoost. Importantly, for CART decision trees, we construct two sets of slowly varying sequences; one using the distance definition defined by [16] and the other using feature importance-based distances. For each model type and each sequence, we graph the AUC and intra-sequence pairwise distance of the models comprising the slowly varying and greedy sequences in Fig. 2. We observe that the test AUC of SVML and greedy sequence grows smaller as we reduce accuracy tolerance, and the absolute difference is small, with a maximum of 1.2% reduction in logistic regression and almost negligible in other model types. However, we observe an obvious advantage in stability measured by pairwise distance, ranging from 5.5% (logistic regression) to 20% (XGBoost).
Figure 2: The AUC and pairwise distances of various machine learning models.
The intra-sequence stability results demonstrate the efficacy of our slowly varying methodology in minimizing the pairwise distance of adjacent models while sacrificing small model accuracy. By setting an accuracy tolerance of $0.01$, for example, we guarantee that the validation AUC of chosen slowly varying models is within $1\%$ of the maximum: a bound that generalizes well experimentally to the test AUC. As expected, the AUC of the models and the pairwise distances for both slowly varying sequences and the greedy sequences tend to increase and decrease over time as they are given more data to be trained on.
Inter-sequence stability.   To test the inter-sequence stability of our slowly varying methodology, we generate multiple different slowly varying XGBoost sequences, each using one more batch of the training data. The sequences were generated by imposing different sequence limits on the optimization problem with an accuracy tolerance of 0.01. We compare inter-sequence stability between the greedy approach, where each end-of-sequence model is selected greedily by performance only, and the SVML approach, which imposes stability by process in Table 1.
Table 1: Inter-Sequence stability between greedy and SVML sequences
Model Name $T_{1-2}$ $T_{2-3}$ $T_{3-4}$ $T_{4-5}$ $T_{5-6}$ $T_{6-7}$ $T_{7-8}$ $T_{8-9}$ $T_{9-10}$ $T_{10-11}$
Greedy 0.476 0.455 0.405 0.363 0.400 0.397 0.353 0.377 0.284 0.255
SVML 0.561 0.418 0.317 0.340 0.294 0.323 0.328 0.233 0.284 0.210
We also select the top 5 most frequently occurring features from the SVML sequence and demonstrate that their respective SHAP importance varies stably across evolving retraining SVML sequences. This qualitatively guarantees that SVML is able to retain analytical insights across different data updates and retraining process.
Figure 3: The absolute average SHAP values (SHAP feature importance) of top-ranking features for the final models of slowly varying sequences at different time intervals, as well as features’ respective average values across all samples
4 Conclusion and Limitations
In practice, retraining may not occur sequentially by incorporating past historical data, but rather by using only the most recent years’ data to avoid data distributional shift. Methodologically, we note that SVML’s deployment must take into consideration the unique computational capacities of the problem studied, and if needed, apply efficient speed-ups to make it scalable. In our existing case study, it takes as much as 2 weeks, with a large portion of the time spent choosing medically-sound candidates to form a candidate pool. Additional challenges, such as when the frequency of data transfer changes (such as from every 24 hours to every 6 hours), may also arise as data updating varies.
In this paper, we propose slowly varying machine learning (SVML), a mixed integer optimization algorithm that trains globally optimal machine learning models across different data batch updates to retain model structural stability. Importantly, our method improves model stability by explicating imposing structural similarities via learning stability by process. We show this algorithm provides significant stability improvements over greedily chosen models focused on performance only, and demonstrate its applicability and versatility with a real-world healthcare case study.

### v2

@@
-Formatting instructions for NeurIPS 2018
+Towards Stable Machine Learning Model Retraining via Slowly Varying Sequences
 Abstract
@@
 1 Introduction
-As more machine learning models are adopted in real-world production pipelines, there is an increasing demand for interpretability to obtain key analytical insights which can build trust in AI-driven decision support systems [1, 2, 3].
-Such model understanding enables developers to evaluate a model’s reliability, which is crucial for high-stake decision spaces such as politics and healthcare [4, 5, 6].
-However, when a machine learning model is updated as new data is incorporated into training, analytical insights often diverge significantly from previous observations [7].
-As users develop trust in models over time based on their past performance and model alignment with personal expertise, when insights suddenly shift due to new data update, it can lead to skepticism and hesitation for adoption .
+The adoption of machine learning models in real-world production pipelines increases the need for interpretability to obtain analytical insights and build trust in AI-driven decision support systems [1, 2, 3].
+Model understanding is crucial for high-stakes decision spaces [4] including healthcare [5, 6] and the social sciences [7]: users develop trust in models over time based on the models’ past performance and alignment with the users’ personal expertise.
+In practice, machine learning models need to frequently be updated as more data becomes available and is incorporated into training; as a result, the model used in production may diverge over time [8].
+Significant shifts in the model’s structure and the resulting analytical insights may may lead to skepticism and hesitation for adoption.
 Consider a predictive model used to estimate a patient’s mortality in a hospital: if the model’s insights drastically change with each new iteration, physicians may find it challenging to explain these variations to their patients and may even create legal difficulties if incidents of error arise.
-Patients rely on doctors’ recommendations, and if the predictions keep fluctuating, it can challenge their confidence in the medical decision-making process [4].
+Patients rely on doctors’ recommendations and,if the predictions fluctuate, it can challenge their confidence in the medical decision-making process [6].
 There are multiple avenues to address challenges faced by retraining in the literature, via continual learning [9], online learning [10], as well as many others [11, 12, 13, 14].
-However, existing methods focus largely on retaining prior performance on the retrained models in the face of data distributional shifts, or conceptual shifts, and thus do not put an emphasis on the fact that analytical insights: such as the decision path of a prediction, or important features selected by a model, be stable across different retraining iterations.
+Existing methods focus largely on retaining prior performance on the retrained models in the face of data distributional shifts, or conceptual shifts, and thus do not put an emphasis on the fact that analytical insights: such as the decision path of a prediction, or important features selected by a model, be stable across different retraining iterations.
 In this work, we demonstrate a direct incorporation of analytical insights into training more stable models and ensure that the output models provide consistent analytical outcomes.
 Additionally, retraining should be distinguished from online learning, which aims at continuously updating data and more importantly, incrementally updating the machine learning model instead of retraining the model from scratch with each data batch update (batch learning).
-One advantage of batch learning is its smaller consumption of computational resources and faster retraining speed since it is done in a periodical instead of a continuous fashion.
+One advantage of batch learning is its smaller consumption of computational resources and faster retraining speed since it is done in a periodical instead of in a continuous fashion.
 Furthermore, such a training strategy allows for a more global solution, as retraining involves using the entire database, whereas online learning, often to ensure faster speed, incrementally incorporates only a portion of the original data, and provides more greedy solutions [15].
@@
 Other works including [18] also developed methods that produce stable feature interactions in random forest-based models.
-Our work follows and extends the work of [16] by developing a unifying framework that learns stable structure across multiple different modeling techniques, including regression, classification trees, and gradient-boosted trees.
+Our work follows and extends the work of [17] by developing a unifying, model-agnostic framework that learns stable structure across multiple different modeling techniques, including regression, classification trees, and gradient-boosted trees.
 1.1 Contributions
-We summarize the main contributions of this paper: (i) we develop a general framework aimed at stabilizing machine learning model structures and formulate it as a mixed-integer optimization-based algorithm for retraining machine learning models that improve existing greedy approaches.
-(ii) we demonstrate a significant boost in model stability with small, computational efficiency, and controlled sacrifice of performance.
-(iii) we show its applicability in an existing production pipeline used in a large US hospital for mortality prediction.
+We summarize the main contributions of this paper: (i) We develop a general framework aimed at stabilizing machine learning model structures upon retraining with new data.
+We propose an exact mixed-integer optimization formulation and a heuristic restricted formulation for the problem, and develop a practical, model-agnostic algorithm for retraining machine learning models.
+(ii) We demonstrate, across a variety of models (including logistic regression, classification trees, and boosted trees), a significant boost in model stability with a small and controllable sacrifice of predictive power and computational efficiency compared to existing greedy approaches.
+(iii) We show our method’s applicability in an existing production pipeline used in a large US hospital for mortality prediction.
 The rest of the paper is organized as follows.
-In Section 2, we outline the problem setting and formulation of the retraining setting, in Section 3, we present numerical experimental results on a real-world case study, in 4, we discuss our limitations and conclude our study.
+In Section 2, we outline the problem setting and develop the proposed retraining methodology; in Section 3, we present numerical experimental results on a real-world case study; in 4, we discuss the limitations and conclude our study.
 2 A Methodology for Retraining Machine Learning Models
-In this section, we introduce the problem setting of the retraining problem, its formulation as a mixed integer problem, and finally discuss the adaptive training strategies.
+In this section, we formalize and formulate the machine learning model retraining problem as an optimization problem, we develop a model-agnostic algorithm for retraining machine learning models, and, finally, we discuss adaptive training strategies.
 2.1 Problem Setting
-Consider an observational dataset composed of triplets $(\boldsymbol{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\boldsymbol{x_{i}}\in\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathbb{R}$ is the target variable, and $t_{i}$ is the time at which the data point $i$ was obtained.
+Consider an observational dataset composed of triplets $(\bm{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\bm{x_{i}}\in\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathbb{R}$ is the target variable, and $t_{i}$ is the time at which data point $i$ was obtained.
 The dataset is accumulated over time, resulting in a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$.
@@
 Minimizing Predictive Error: The first requirement is to ensure that each model $f_{b}$ in the sequence $f_{1},\ldots,f_{B}$ has a low out-of-sample predictive error.
-Formally, denoting by $\mathcal{D}_{b}$ the distribution of the training data in $D_{b}$, we seek to minimize the expected error for data points $(\boldsymbol{x},y)\sim\mathcal{D}_{b}$:
-$\min_{f_{1},\ldots,f_{B}}\sum_{b=1}^{B}\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}_{b}}[\ell(f_{b}(\boldsymbol{x}),y)],$ (1)
-where the loss function $\ell(\cdot,\cdot)$ quantifies the prediction error when the predicted value is $f_{b}(\boldsymbol{x})$ and the true value is $y$.
+Formally, denoting by $\mathcal{D}_{b}$ the distribution of the training data in $D_{b}$, we seek to minimize the expected error for data points $(\bm{x},y)\sim\mathcal{D}_{b}$:
+$\min_{f_{1},\ldots,f_{B}}\sum_{b=1}^{B}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}_{b}}[\ell(f_{b}(\bm{x}),y)],$ (1)
+where the loss function $\ell(\cdot,\cdot)$ quantifies the prediction error when the predicted value is $f_{b}(\bm{x})$ and the true value is $y$.
 Ensuring Model Stability: The second requirement is to ensure that consecutive model structures do not vary too significantly, maintaining stability over time.
@@
 $\begin{split}\min_{f_{1},\ldots,f_{B}}&\quad\sum_{b=1}^{B-1}d(f_{b},f_{b+1})\\ \text{s.t.}&\quad L(f_{b},D_{b})\leq(1+\alpha)\cdot\min_{f}L(f,D_{b})\quad\forall b\in[B].\end{split}$ (4)
-Depending on the type of the model used, solving Problem (4) can be challenging, as we need to consider either an infinite number (i.e.
-neural network) or an exponentially growing number (with respective to feature size, i.e.
-decision tree) of possible functions $f_{b}$.
-This is particularly true for models that are commonly used in high-stakes healthcare applications, e.g., decision trees and variants thereof.
+Depending on the type of the model used, solving Problem (4) can be challenging: the number of possible functions $f_{b}$ may grow exponentially with the number of features.
+This is particularly true for models that are commonly used in high-stakes healthcare applications, including decision trees and variants thereof: considering $p$ binary features and a binary classification task, then there are $2^{2^{p}}$ different ways of labeling all instances, each corresponding to a different underlying boolean function that can be represented as a decision tree.
 Thus, we next develop a restricted formulation that can tractably accommodate any type of machine learning model.
@@
 Structural Distances: This approach measures the distance by directly comparing the models’ structures. We provide the following concrete examples:
-  • For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\beta}^{i}-\boldsymbol{\beta}^{j}\|_{2}^{2}$, where $\boldsymbol{\beta}^{i}$ and $\boldsymbol{\beta}^{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively.
+  • For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\bm{\beta}^{i}-\bm{\beta}^{j}\|_{2}^{2}$, where $\bm{\beta}^{i}$ and $\bm{\beta}^{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively.
   • For decision tree-based models, we use the decision tree distance metric proposed by [17] that relies on optimally matching the paths in the two trees $\mathbb{T}_{1}$ and $\mathbb{T}_{2}$, denoted as $\mathcal{P}(\mathbb{T}_{1})$ and $\mathcal{P}(\mathbb{T}_{2})$.
-The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account a path weight $w(p)$, defined as the differences in feature ranges in the two paths as well as the class labels associated with them.
+The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account the differences in feature ranges in the two paths as well as the class labels associated with them.
 Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem.
-This distance is defined as an integer linear optimization problem, where given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{1}$ is matched with path $q$ in $\left.\mathbb{T}_{2}\right)$ and $x_{p}=$ $\mathbb{I}$ (path $p$ in $\mathbb{T}_{1}$ is left unmatched), we formulate it as below:
-    $d\left(\mathbb{T}_{1},\mathbb{T}_{2}\right)=\min_{\boldsymbol{x}}$ $\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}\sum_{q\in\mathcal{P}\left(\mathbb{T}_{2}\right)}d(p,q)x_{pq}+\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}w(p)x_{p}$
-    s.t. $\sum_{q\in\mathcal{P}\left(\mathbb{T}_{2}\right)}x_{pq}+x_{p}=1,\quad\forall p\in\mathcal{P}\left(\mathbb{T}_{1}\right)$
-    $\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}x_{pq}=1,\quad\forall q\in\mathcal{P}\left(\mathbb{T}_{2}\right)$
-    $x_{pq}\in\{0,1\},\quad x_{p}\in\{0,1\},\quad\forall p\in\mathcal{P}\left(\mathbb{T}_{1}\right),\quad\forall q\in\mathcal{P}\left(\mathbb{T}_{2}\right)$
+Formally, assuming without loss of generality that $\mathbbm{T}_{1}$ consists of a larger number of paths, and given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{1}$ is matched with path $q$ in $\left.\mathbb{T}_{2}\right)$, the distance is defined as the following integer linear optimization problem:
+    $\begin{split}d(\mathbbm{T}_{1},\mathbbm{T}_{2})\ =\ \min_{\bm{x}}&\quad\sum_{p\in\mathcal{P}(\mathbb{T}_{1})}\sum_{q\in\mathcal{P}(\mathbb{T}_{2})}d(p,q)x_{pq}\\ \text{s.t.}&\quad\sum_{q\in\mathcal{P}(\mathbb{T}_{2})}x_{pq}=1,\quad\forall p\in\mathcal{P}(\mathbb{T}_{1})\\ &\quad\sum_{p\in\mathcal{P}(\mathbb{T}_{1})}x_{pq}\geq 1,\quad\forall q\in\mathcal{P}(\mathbb{T}_{2})\\ &\quad x_{pq}\in\{0,1\},\quad\forall p\in\mathcal{P}(\mathbb{T}_{1})\ \forall q\in\mathcal{P}(\mathbb{T}_{2}).\end{split}$ (6)
 Feature Importance-Based Distances: This approach measures the distance by comparing the models based on their respective feature importances.
-In particular, denoting by $\boldsymbol{\phi}^{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\phi}^{f_{i}}-\boldsymbol{\phi}^{f_{j}}\|_{2}^{2}$.
+In particular, denoting by $\bm{\phi}^{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\bm{\phi}^{f_{i}}-\bm{\phi}^{f_{j}}\|_{2}^{2}$.
 The advantage of using feature importance-based distances is its versatility to be applied to both interpretable (decision-tree and logistic regression), as well as black-box (gradient boosted tree) models.